OpenStudy (anonymous):

2.25(4x-4) = -2 + 10x + 12 =

OpenStudy (gorv):

just take 2.25 inside and distribute

OpenStudy (gorv):

2.25*4=9

OpenStudy (gorv):

can u do it further

OpenStudy (anonymous):

@gorv let @Breana86 participate too, you're going too fast.

OpenStudy (gorv):

yeah i m doing that also

OpenStudy (anonymous):

Yes thank you so much gorv, that was awesome, helping my 14 year old brother and we are having lots of trouble with this part

OpenStudy (anonymous):

Nope still cant get any futher @gorv

OpenStudy (anonymous):

Thats what we dont get, why do we do that?

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

we need to find x so the element we need to find we keep that on one side move other things on other side

OpenStudy (anonymous):

So do we add 10 to 9x or 9? Thank you for your help @gorv

OpenStudy (gorv):

we add or subtract similar term only

OpenStudy (gorv):

9x and 10 are not similar

OpenStudy (anonymous):

so 9 and 10 are? @gorve

OpenStudy (gorv):

yeah

OpenStudy (anonymous):

So now we have 9x - 19 = 10x, is that correct @gorv

OpenStudy (anonymous):

We dont know what to do now?

OpenStudy (anonymous):

Where did you get -9 from? @gorv

OpenStudy (gorv):

9x-9+10=9x+10-9 now can do it

OpenStudy (anonymous):

I am so sorry, we just are not getting it! Now we have 9x - 19 = 9x 1 @gorv

OpenStudy (anonymous):

Sorry we are just not getting it, the back of the book says its -19 but we still dont know how to come up with that @gorv

OpenStudy (gorv):

9x-9=10x+10 we need to subtacrt10 from both side

OpenStudy (gorv):

u r right 9x-19=10x

OpenStudy (gorv):

now subtract 9x from both side

OpenStudy (gorv):

@Breana86

OpenStudy (anonymous):

so now we have 1x = -19 @gorv

OpenStudy (gorv):

yeah thatzz it

OpenStudy (gorv):

1x=x x=-19

OpenStudy (anonymous):

YAYAYYAYAYAYAYAAY Thank you!! Now on to the next! @gorv

OpenStudy (gorv):

well post new Q ..i have to go now loll.....someone will surely help u :)